# Bayesian Posteriors with Stellar Population Synthesis on GPUs

### v1

@@ Abstract
Models of Stellar Population Synthesis (SPS) provide a predictive framework for the spectral energy distribution (SED) of a galaxy. SPS predictions can be computationally intensive, creating a bottleneck for attempts to infer the physical properties of large populations of individual galaxies from their SEDs and photometry; these computational challenges are especially daunting for near-future cosmology surveys that will measure the photometry of billions of galaxies. In this paper, we explore a range of computational techniques aimed at accelerating SPS predictions of galaxy photometry using the JAX library to target GPUs. We study a particularly advantageous approximation to the calculation of galaxy photometry that speeds up the computation by a factor of 50 relative to the exact calculation. We introduce a novel technique for incorporating burstiness into models of galaxy star formation history that captures very short-timescale fluctuations with negligible increase in computation time. We study the performance of Hamiltonian Monte Carlo (HMC) algorithms in which individual chains are parallelized across independent GPU threads, finding that our pipeline can carry out Bayesian inference at a rate of approximately $1,000$ galaxy posteriors per GPU-minute. Our results provide an update to standard benchmarks in the literature on the computational demands of SPS inference; our publicly available code enables previously-impractical Bayesian analyses of large galaxy samples, and includes several standalone modules that could be adopted to speedup existing SPS pipelines.
@@
Cosmology: large-scale structure of Universe
@@ 1 Introduction
The prevailing theoretical framework for predicting the spectral energy distribution (SED) of a galaxy from its fundamental physical properties is stellar population synthesis (SPS, Conroy, 2013). The concepts and defining equations of SPS have been in place for decades (e.g., Tinsley, 1978; Bruzual A., 1983; Arimoto & Yoshii, 1987; Buzzoni, 1989; Bruzual A. & Charlot, 1993; Worthey, 1994; Maraston, 1998; Leitherer et al., 1999), and there are numerous SPS libraries that are publicly available (e.g., Fioc & Rocca-Volmerange, 1999; Bruzual & Charlot, 2003; Le Borgne et al., 2004; Maraston, 2005; Conroy et al., 2009; Eldridge et al., 2017; Bradley et al., 2020; Johnson et al., 2021).
@@
One of the most common forms of SPS analysis is fitting the observed SED or photometry of a galaxy with a physical model (e.g., Sawicki & Yee, 1998; Brinchmann & Ellis, 2000; Salim et al., 2007; Kriek et al., 2009). Monte Carlo Markov Chains (MCMCs) are the standard technique for Bayesian uncertainty quantification in such fits, and SPS inference with MCMC can be computationally costly. As a case study for the computational demands of SPS inference, consider Leja et al. (2019b), which used a 14-parameter SPS model implemented in Prospector (Johnson et al., 2021) to analyze 58,461 galaxies from the 3D-HST survey (Momcheva et al., 2016); the authors report using $\sim 10^{5}-10^{6}$ likelihood evaluations for each galaxy’s MCMC chain to converge, and an average runtime of 25 CPU hours per galaxy, resulting in a total of 1.5 million CPU hours for the analysis. [^1] Comparable per-galaxy MCMC runtimes with Prospector are also reported in Tacchella et al. (2022). For cosmological datasets with $10^{7}-10^{9}$ galaxies, naively scaling traditional SPS inference techniques would require hundreds of millions to billions of CPU-hours; these demands are especially daunting considering that obtaining robust results commonly requires many MCMCs per galaxy, for example when running tests for the effect of different spectral libraries in the analysis.
@@
There are essentially two classes of approaches to addressing this computational challenge: speed up the computation of the SPS prediction for the likelihood, or use some alternative to running a traditional MCMC for each galaxy in the sample. Using an emulator of the SPS predictions is a natural way to speed up the cost of evaluating the likelihood. In Hahn et al. (2023a), a neural network emulator of their PROVABGS model was used to derive posteriors on galaxies in DESI BGS (Hahn et al., 2023b). The authors found that their SPS model evaluates in 340 ms compared to 2.9 ms with the emulator, and that with their emulator they achieve MCMC runtimes of $\sim 5$ minutes per galaxy. Large performance gains from neural emulation are also reported in Alsing et al. (2020), in which the authors use a training set of 2 million synthetic galaxy SEDs and photometry to train an emulator of SPS predictions by Prospector, finding speedup factors of $10^{3}-10^{4}$.
@@
As an example of the second class of approach, in Hahn & Melchior (2022) the authors developed a neural density estimator of individual galaxy posteriors, SEDFlow. To train SEDflow, they use the PROVABGS model to generate MCMC posteriors of $\sim 10^{6}$ synthetic galaxies; once trained, SEDflow can infer the posterior of each new galaxy in $\sim 1$ CPU-second. When emulating per-galaxy posteriors, the computational demands become driven by size of the training data required to train an accurate network, rather than by the size of the observational dataset.
@@
An alternative, but closely related, approach to per-galaxy MCMCs is population-level modeling, which adopts a different strategy to analyze the galaxy population. In population-level modeling, some parametric model (typically a neural network) is used to describe the PDF of physical galaxy properties, and the parameters of the model are tuned so that the galaxy population generated by the PDF presents summary statistics that are in close agreement with the corresponding observations. In Li et al. (2024), the authors use a population-level model, POPSED, to recover various summary statistics of the population of galaxies observed in the GAMA survey (Driver et al., 2011). Another population-level model, Pop-Cosmos (Alsing et al., 2024), was used to capture the statistical distribution of galaxy properties modeled by Prospector, and thereby place constraints on several key scaling relations of the physics of galaxies in the COSMOS-20 dataset.
@@
In the present work, we study three computational techniques for improving the efficiency of SPS inference. First, we explore a method for approximating the effect of dust attenuation on the photometry calculation. Second, we evaluate the computational gains derived from evaluating the likelihood on a GPU using the DSPS library (Hearin et al., 2023). Finally, we study the computational benefits of leveraging gradient information in the inference algorithm. Broadly, we find that with these techniques used in tandem, we reduce the computation time to derive per-galaxy posteriors by factors of 100-1000 relative to running conventional MCMCs on a CPU.
@@ 2 SPS model overview
×$\times P_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|z,\theta_{\rm gal})\;.$ (1)
In Equation (2), $M_{\star}$ is the total stellar mass formed by redshift $z_{\rm obs},$ $F_{\rm att}(\lambda)$ is the attenuation function that encodes the absorption of light by dust within the host galaxy, and $P_{\rm SSP}(\tau_{\rm age},Z)$ is the probability distribution of stars in the galaxy as a function of age and metallicity. In typical SPS analyses, a particular choice is made in advance for the IMF, isochrones, and stellar SED library, and thereafter the SSP SEDs $L_{\rm SSP}(\lambda|\tau_{\rm age},Z)$, are held fixed throughout the computations; we adopt SSP SEDs from MIST (Choi et al., 2016) throughout this work.
@@ 3 Computing approximate photometry
[…]
@@
×$\times\int d\lambda\;T(\lambda|z)L_{\rm SSP}(\lambda|\tau_{\rm age}^{i},Z^{j})F_{\rm att}(\lambda|\theta_{\rm gal}).$ (4)
@@
[…]
@@
×$\qquad\times\sum_{i,j}P_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|\theta_{\rm gal})\cdot c_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|z)$
@@
[…]
The quantities $c_{\rm SSP}$ can be calculated in advance of the analysis at each point on the SSP grid, either at a fixed $z$ if the redshift is known, or on a lookup table, so that calculating the photometry of a model galaxy only requires computing $P_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|\theta_{\rm gal}),$ and then taking $P_{\rm SSP}$-weighted sums of the grid of $c_{\rm SSP}.$ Variations on this approximation are used by several SSP libraries, for example in CIGALE (Burgarella et al., 2005; Noll et al., 2009; Boquien et al., 2019) and the SEDITION code included in UniverseMachine (Behroozi et al., 2019, 2020), and in this paper we provide systematic testing of its accuracy and performance benefits in the context of SPS inference.
@@ 4 GPU-Parallel SPS Inference
In this section, we study the computational problem of deriving Bayesian posteriors, $\mathcal{P}(\theta_{\rm gal}|D),$ on the parameters $\theta_{\rm gal}$ of an individual galaxy constrained by observations of its photometry, $D.$ As discussed in Section 1, runtimes of $\sim 25$ CPU-hour per galaxy are often quoted in the literature for SPS models using variants of MCMC methods to estimate the posterior. Here we revisit this computational benchmark using Hamiltonian Monte Carlo (HMC, Duane et al., 1987; Neal, 2011), an alternative Bayesian algorithm that leverages gradient information, and using SPS computations that have been optimized for GPU computing resources.
@@
Using the SPS model outlined in Section 2, we generate photometry for hundreds of thousands of synthetic galaxies by randomly sampling $\theta_{\rm gal}$ from a broad prior (see Appendix C for detailed information about the prior distribution we adopted). We treat this photometry as target data, and we run an HMC chain for each synthetic galaxy using the implementation in Blackjax, a library of samplers written in JAX that can be run on both CPU and GPU machines. We run each chain for a fixed number of steps and check for convergence according to the Gelman-Rubin (Gelman & Rubin, 1992) statistic, which compares the variance between and within chains to determine if a run has converged to a stationary state; we adopt $\hat{R}<1.2$ as a convergence threshold for each chain.
@@
In Figure 2 we show the runtime for our Bayesian inference pipeline. The horizontal axis shows how many galaxies are in the sample for which posteriors were derived; the vertical axis shows the wall-clock time in node minutes; we show computational performance for three different exercises: HMC run on GPUs and CPUs, as well as an alternative gradient-based posterior sampler implemented in Blackjax, the No-U-Turn Sampler (NUTS, Hoffman & Gelman, 2014), on GPUs. We run each sampler for $N_{\rm steps}=3000$ steps, with an adaptation phase of $100$ steps, using the exact same target data and sampler settings. We selected $N_{\rm steps}$ to strike a balance between having a high fraction of chains that fully converge, and having a fast runtime. For $N_{\rm steps}=3000$, more than $85\%$ of the fits have fully converged for both HMC and NUTS; the remaining fits are typically close to convergence, and require an additional $\sim N_{\rm steps};$ a small portion of outliers typically requires adjusting the starting point of the chain.
@@
The most striking feature of Figure 2 is the performance of HMC on GPUs: there is virtually no increase in wall-clock time with $N_{\rm chains}\lesssim 2000.$ This reflects both the nature of the GPU architecture as well as the design of our pipeline. GPU computations achieve especially high performance on problems where the same work runs in parallel across many independent GPU threads. For the case of the Blackjax implementation of HMC, it is straightforward to parallelize the computation so that each chain is run on an independent GPU thread; since the HMC algorithm does not require iterative calculations in between steps, then each chain requires the same amount of work, and so the chains proceed in lock-step with each other for the pre-determined number of samples. With these considerations in mind, our strategy to optimize these computations is to fit as many independent chains as can fit on the GPU memory; once the available memory is saturated, the runtime begins to rise sharply, which for our benchmark we find occurs for $N_{\rm chains}\gtrsim 10^{4}.$
@@ 5 Conclusion
We thus expect that AI-based methods will continue to play an important role in SPS analyses of large galaxy samples. There are numerous such applications for which the pipeline we have presented here could provide useful support. For example, high-accuracy AI emulators of photometry and SEDs require large samples of training data, and DSPS is able to generate such training data in a fraction of the runtime of comparable SPS libraries. Techniques such as amortized Neural Posterior Estimation (e.g., SEDflow, Hahn & Melchior, 2022) require large samples of posteriors for training, and the computational gains of our GPU-parallel pipeline makes it far more practical to create such data. Recent work in cosmological inference has demonstrated the benefits of likelihood emulation (To et al., 2023), and the techniques presented here could provide training sets of both likelihood evaluations as well as gradients of the likelihood. Recent advances in Simulation Based Inference (Zeghal et al., 2022) have shown considerable improvements in sampling efficiency when leveraging the availability of gradient information. We have developed our GPU-parallel pipeline with these and other applications in mind for future work.

### v2

@@ Abstract
Models of Stellar Population Synthesis (SPS) provide a predictive framework for the spectral energy distribution (SED) of a galaxy. SPS predictions can be computationally intensive, creating a bottleneck for attempts to infer the physical properties of large populations of individual galaxies from their SEDs and photometry; these computational challenges are especially daunting for near-future cosmology surveys that will measure the photometry of billions of galaxies. In this paper, we explore a range of computational techniques aimed at accelerating SPS predictions of galaxy photometry using the JAX library to target GPUs. We study a particularly advantageous approximation to the calculation of galaxy photometry that speeds up the computation by a factor of 50 relative to the exact calculation. We introduce a novel technique for incorporating burstiness into models of galaxy star formation history that captures very short-timescale fluctuations with negligible increase in computation time. We study the performance of Hamiltonian Monte Carlo (HMC) algorithms in which individual chains are parallelized across independent GPU threads, finding that our pipeline can carry out Bayesian inference at a rate of approximately $1000$ galaxy posteriors per minute on a single GPU. Our results provide an update to standard benchmarks in the literature on the computational demands of SPS inference; our publicly available code enables previously-impractical Bayesian analyses of large galaxy samples, and includes several standalone modules that could be adopted to speedup existing SPS pipelines.
@@
Cosmology: large-scale structure of Universe, Galaxies: fundamental parameters
@@ 1 Introduction
The prevailing theoretical framework for predicting the spectral energy distribution (SED) of a galaxy from its fundamental physical properties is stellar population synthesis (SPS, conroy13_sps_aa_review). The concepts and defining equations of SPS have been in place for decades (e.g., Tinsley1978_sps; Bruzual1983_sps; Arimoto1987_sps; Buzzoni1989_sps; bruzual_charlot_1993; worthey_1994_age_metal_degeneracy; maraston_1998; Leitherer_starburst99), and there are numerous SPS libraries that are publicly available (e.g., fioc_rocca_volmerange99_pegase; Bruzual2003; leborgne_etal04; maraston05; conroy_gunn_fsps1; eldridge_etal17; larry_bradley_2020_photutils_v1p0; johnson_etal21_prospector).
@@
One of the most common forms of SPS analysis is fitting the observed SED or photometry of a galaxy with a physical model (e.g., sawicki_yee_1998; brinchmann_ellis_2000; salim_etal07_sdss_sfrs; kriek_etal09). Monte Carlo Markov Chains (MCMCs) are the standard technique for Bayesian uncertainty quantification in such fits, and SPS inference with MCMC can be computationally costly. As a case study for the computational demands of SPS inference, consider leja_3dhst_2019a, which used a 14-parameter SPS model implemented in Prospector (johnson_etal21_prospector) to analyze 58,461 galaxies from the 3D-HST survey (momcheva_etal16_3dhst); the authors report using $\sim 10^{5}-10^{6}$ likelihood evaluations for each galaxy’s MCMC chain to converge, and an average runtime of 25 CPU hours per galaxy, resulting in a total of 1.5 million CPU hours for the analysis. [^1] Comparable per-galaxy MCMC runtimes with Prospector are also reported in tacchella_etal22. For cosmological datasets with $10^{7}-10^{9}$ galaxies, naively scaling traditional SPS inference techniques would require hundreds of millions to billions of CPU-hours; these demands are especially daunting considering that obtaining robust results commonly requires many MCMCs per galaxy, for example when running tests for the effect of different spectral libraries in the analysis.
@@
There are essentially two classes of approaches to addressing this computational challenge: speed up the computation of the SPS prediction for the likelihood, or use some alternative to running a traditional MCMC for each galaxy in the sample. Using an emulator of the SPS predictions is a natural way to speed up the cost of evaluating the likelihood. In hahn_etal23_PROVABGS, a neural network emulator of their PROVABGS model was used to derive posteriors on galaxies in DESI BGS (hahn_etal23_desi_bgs). The authors found that their SPS model evaluates in 340 ms compared to 2.9 ms with the emulator, and that with their emulator they achieve MCMC runtimes of $\sim 5$ minutes per galaxy. Large performance gains from neural emulation are also reported in alsing_etal20_speculator, in which the authors use a training set of 2 million synthetic galaxy SEDs and photometry to train an emulator of SPS predictions by Prospector, which they call Speculator, finding speedup factors of $10^{3}-10^{4}$.
@@
As an example of the second class of approach, in hahn_melchior_2022_accel_bayes_sed the authors developed a neural density estimator of individual galaxy posteriors, SEDFlow. To train SEDflow, they use the PROVABGS model to generate MCMC posteriors of $\sim 10^{6}$ synthetic galaxies; once trained, SEDflow can infer the posterior of each new galaxy in $\sim 1$ CPU-second. When emulating per-galaxy posteriors, the computational demands become driven by size of the training data required to train an accurate network, rather than by the size of the observational dataset.
@@
An alternative, but closely related, approach to per-galaxy MCMCs is population-level modeling, which adopts a different strategy to analyze the galaxy population. In population-level modeling, some parametric model (typically a neural network) is used to describe the PDF of physical galaxy properties, and the parameters of the model are tuned so that the galaxy population generated by the PDF presents summary statistics that are in close agreement with the corresponding observations. In li_etal24_popsed, the authors use a population-level model, POPSED, to recover various summary statistics of the population of galaxies observed in the GAMA survey (driver_etal11_gama_survey). Another population-level model, Pop-Cosmos (alsing_etal24_pop_cosmos), was used to capture the statistical distribution of galaxy properties modeled by Prospector, and thereby place constraints on several key scaling relations of the physics of galaxies in the COSMOS-20 dataset. In subsequent studies, in Thorp_etal24_popcosmos; Thorp_etal25_popcosmos Pop-Cosmos was used to estimate properties and redshifts for individual galaxies. The authors quote runtimes of 15 GPU-second per galaxy using the pop-cosmos prior, and 0.6 GPU-second per galaxy under the Prospector prior, a significant milestone in the development of fast and accurate SPS parameter posterior inference.
@@
In the present work, we study three computational techniques for improving the efficiency of SPS inference. First, we explore a method for approximating the effect of dust attenuation on the photometry calculation. Second, we evaluate the computational gains derived from evaluating the likelihood on a GPU using the DSPS library (Hearin2023DSPS). Finally, we study the computational benefits of leveraging gradient information in the inference algorithm. Broadly, we find that with these techniques used in tandem, we reduce the computation time to derive per-galaxy posteriors by factors of 100-1000 relative to running conventional MCMCs on a CPU.
@@ 2 SPS model overview
$\times P_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|z,\theta_{\rm gal})\;.$ (1)
In Equation (2), $M_{\star}$ is the total stellar mass formed by redshift $z_{\rm obs},$ $F_{\rm att}(\lambda)$ is the attenuation function that encodes the absorption of light by dust within the host galaxy, and $P_{\rm SSP}(\tau_{\rm age},Z)$ is the probability distribution of stars in the galaxy as a function of age and metallicity. In typical SPS analyses, a particular choice is made in advance for the IMF, isochrones, and stellar SED library, and thereafter the SSP SEDs $L_{\rm SSP}(\lambda|\tau_{\rm age},Z)$, are held fixed throughout the computations; we adopt SSP SEDs from MIST (choi_etal16_mist) throughout this work.
@@ 3 Computing approximate photometry
$c_{\rm gal}(\theta_{\rm gal},z)\equiv\int d\lambda\;T(\lambda|z)L_{\rm gal}(\lambda|\theta_{\rm gal},z)\;,$ (3)
@@
$\times\int d\lambda\;T(\lambda|z)L_{\rm SSP}(\lambda|\tau_{\rm age}^{i},Z^{j})F_{\rm att}(\lambda|\theta_{\rm gal}).$ (4)
@@
$\lambda_{\rm eff}(z)\equiv\frac{\int d\lambda\;\lambda\cdot T(\lambda|z)}{\int d\lambda\;T(\lambda|z)}\;.$ (5)
@@
$\qquad\times\sum_{i,j}P_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|\theta_{\rm gal})\cdot c_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|z)$
@@
$c_{\rm SSP}(\tau_{\rm gal}^{i},Z^{j}|z)\equiv\int d\lambda\;T(\lambda|z)L_{\rm SSP}(\lambda|\tau_{\rm age}^{i},Z^{j})\;.$ (7)
The quantities $c_{\rm SSP}$ can be calculated in advance of the analysis at each point on the SSP grid, either at a fixed $z$ if the redshift is known, or on a lookup table, so that calculating the photometry of a model galaxy only requires computing $P_{\rm SSP}(\tau_{\rm age}^{i},Z^{j}|\theta_{\rm gal}),$ and then taking $P_{\rm SSP}$-weighted sums of the grid of $c_{\rm SSP}.$ Variations on this approximation are used by several SSP libraries, for example in CIGALE (burgarella_etal05_cigale; noll_etal09_cigale; boquien_etal19_cigale) and the SEDITION code included in UniverseMachine (behroozi_etal19; behroozi_etal20_umachine_jwst), and in this paper we provide systematic testing of its accuracy and performance benefits in the context of SPS inference.
@@ 4 GPU-Parallel SPS Inference
In this section, we study the computational problem of deriving Bayesian posteriors, $\mathcal{P}(\theta_{\rm gal}|D),$ on the parameters $\theta_{\rm gal}$ of an individual galaxy constrained by observations of its photometry, $D.$ As discussed in Section 1, runtimes of $\sim 25$ CPU-hour per galaxy are often quoted in the literature for SPS models using variants of MCMC methods to estimate the posterior. Here we revisit this computational benchmark using Hamiltonian Monte Carlo (HMC, duan_etal87_hmc; neal_2011_hmc_chapter), an alternative Bayesian algorithm that leverages gradient information, and using SPS computations that have been optimized for GPU computing resources. We note that the model we utilized for the benchmarks is not as complex as some SPS models that were employed in the studies whose quoted runtimes we mentioned above. The DSPS library underlying this work is not as full-featured as other, more mature SPS libraries in the field, and in future work we aim to incorporate more realistically complex models into our pipeline. We expect the general trends in computational speedup we present here to continue to hold, since the main reasons for our performance gains are the GPU capability of our code, combined with the ability to leverage gradient-based algorithms, both of which will remain true due to our JAX-based implementation.
@@
Using the SPS model outlined in Section 2, we generate photometry for hundreds of thousands of synthetic galaxies by randomly sampling $\theta_{\rm gal}$ from a broad prior (see Appendix C for detailed information about the prior distribution we adopted). We treat this photometry as target data, and we run an HMC chain for each synthetic galaxy using the implementation in Blackjax, a library of samplers written in JAX that can be run on both CPU and GPU machines. We run each chain for a fixed number of steps and check for convergence according to the Gelman-Rubin (GelmanRubinRhat) statistic, which compares the variance between and within chains to determine if a run has converged to a stationary state; we adopt $\hat{R}<1.2$ as a convergence threshold for each chain.
@@
In Figure 2 we show the runtime for our Bayesian inference pipeline. The horizontal axis shows how many galaxies are in the sample for which posteriors were derived; the vertical axis shows the wall-clock time in node minutes; we show computational performance for three different exercises: HMC run on GPUs and CPUs, as well as an alternative gradient-based posterior sampler implemented in Blackjax, the No-U-Turn Sampler (NUTS, hoffman_gelman_2014_nuts), on GPUs. We run each sampler for $N_{\rm steps}=3000$ steps, with an adaptation phase of $100$ steps, using the exact same target data and sampler settings. We selected $N_{\rm steps}$ to strike a balance between having a high fraction of chains that fully converge, and having a fast runtime. For $N_{\rm steps}=3000$, more than $85\%$ of the fits have fully converged for both HMC and NUTS; the remaining fits are typically close to convergence, and require an additional $\sim N_{\rm steps};$ a small portion of outliers typically requires adjusting the starting point of the chain.
@@
The most striking feature of Figure 2 is the performance of HMC on GPUs: there is virtually no increase in wall-clock time with $N_{\rm chains}\lesssim 2000.$ This reflects both the nature of the GPU architecture as well as the design of our pipeline. GPU computations achieve especially high performance on problems where the same work runs in parallel across many independent GPU threads. For the case of the Blackjax implementation of HMC, it is straightforward to parallelize the computation so that each chain is run on an independent GPU thread; since the HMC algorithm does not require iterative calculations in between steps, then each chain requires the same amount of work, and so the chains proceed in lock-step with each other for the pre-determined number of samples. With these considerations in mind, our strategy to optimize these computations is to fit as many independent chains as can fit in the GPU memory; once the available memory is saturated, the runtime begins to rise sharply, which for our benchmark we find occurs for $N_{\rm chains}\gtrsim 10^{4}$. For reference, on the Swing machine we used at Argonne [^2] https://www.lcrc.anl.gov/systems, a single GPU has 4040GB of memory; on Amazon Web Services (AWS), one typically uses instances with 1616GB memory, or up to 8080GB for the more powerful systems. Doubling the GPU memory doubles the number of galaxies that can be processed at once on a single GPU. For galaxy samples too large to fit in the memory of a single GPU, it is straightforward to divide the sample into subsets, and process each subset on a different GPU in parallel with MPI.
@@ 5 Conclusion
We thus expect that AI-based methods will continue to play an important role in SPS analyses of large galaxy samples. There are numerous such applications for which the pipeline we have presented here could provide useful support. For example, high-accuracy AI emulators of photometry and SEDs require large samples of training data, and DSPS is able to generate such training data in a fraction of the runtime of comparable SPS libraries. Techniques such as amortized Neural Posterior Estimation (e.g., SEDflow, hahn_melchior_2022_accel_bayes_sed) require large samples of posteriors for training, and the computational gains of our GPU-parallel pipeline makes it far more practical to create such data. Recent work in cosmological inference has demonstrated the benefits of likelihood emulation (to_etal23_linna), and the techniques presented here could provide training sets of both likelihood evaluations as well as gradients of the likelihood. Recent advances in Simulation Based Inference (zeghal_etal_22_sbi_gradients) have shown considerable improvements in sampling efficiency when leveraging the availability of gradient information. We have developed our GPU-parallel pipeline with these and other applications in mind for future work.